OpenStudy (anonymous):

|4-1/2x|<=6 *sn: there is a line under the < symbol

OpenStudy (amistre64):

4-1/2x <=6 and -4+1/2x <=6 -1/2x <= 2 1/2x <=10 x >= -4 x <= 20 is what I get if I kept it straight

OpenStudy (anonymous):

thank you...you are a pro at this lol

OpenStudy (anonymous):

i wish i could give you more medals

OpenStudy (amistre64):

lol..... lots a memory dust to sift thru ;)

OpenStudy (amistre64):

medals are a new fad in here lol... I just like answering the math really :)

OpenStudy (anonymous):

lol well thats good too :)

OpenStudy (anonymous):

It's a bit easier IMO to evaluate it as a single relation. \[|4-1/2x| \le 6\] \[\implies -6 \le 4-(1/2)x \le 6\] \[\implies -10 \le -(1/2)x \le 2\] \[\implies 20 \ge x \ge -4\]

OpenStudy (amistre64):

everytime a try a single relation, I mess it up.... ;)

OpenStudy (anonymous):

Oh, I find it much easier to keep straight that way ;p